OpenStudy (anonymous):

can someone help me to find the factor for this 8r^3 - 64t^3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

think about difference of cubes (x^3 - y^3) and that 64 is 4^3 and 8 is 2^3

OpenStudy (skullpatrol):

$$8r^3 - 64t^3$$

OpenStudy (skullpatrol):

$$x^3-y^3=(x-y)(x^2+xy+y^2)$$

OpenStudy (anonymous):

I am totally lost

OpenStudy (anonymous):

8r^3 = (2r)^3 right? and 64t^3=(4t)^3 so lets call 2r = x and 4t = y that would give me x^3 - y^3 for ur initial expression right? if u plug in skull formula and after change x and y to their originals u'll get the answer